OpenStudy (anonymous):

The sum of two numbers is 110. Their difference is 28. What are the two numbers. Write them like this example, with no spaces: 15and16.

OpenStudy (anonymous):

x+y=110 x-y=28 add the 2 equations: 2x=138 x=69 y=x-28=41 41 and 69